“I am four times as old as you were when I was the same age as you are now. I am forty years old, how old are you?”
Answer by ptaylor(1894)

Let x=your present age
And 40=my present age
Some "y" years ago I was the same age as you are now, so:
40-y=x
-y=x-40
y=40-x
So 40-x years ago I was the same age as you are now
And 40-x years ago, you were x-(40-x)=2x-40 years old.
So, our equation to solve is:
40=4*(2x-40)
40=8x-160
8x=200
x=25----------------your present age
CK
15 years ago, I was the same age as your are now (40-15=25)
15 years ago, you were 25-15=10
40=4*10
Hope this helps---ptaylor